Please help me solve this question. ------------- o -- -- (1 point) -6 Fill in the blanks to describe the given piecewise graph. Hint: In each row, the first bla
Answer #1

From the graph, we observe that

For -5 < x < 0

The graph passes through the points (-5, -1) and (0, -3)

Therefore, the slope m = (-3 - (-1)) / (0 - (- 5)) = - 2 / 5

Equation of the line is

y - (-1) = - (2 / 5) (x - (-5))

=> y + 1 = (-2/5)(x + 5)

=> y + 1 = (-2/5)x - 2

=> y = (-2/5)x - 3

Therefore, f(x) = (-2/5)x - 3

For 0 < x < 5

The graph passes through the points (5, -1) and (0, -3)

Therefore, the slope m = (-3 - (-1)) / (0 - (5)) = 2 / 5

Equation of the line is

y - (-1) = (2 / 5) (x - 5)

=> y + 1 = (2/5)(x - 5)

=> y + 1 = (2/5)x - 2

=> y = (2/5)x - 3

Therefore, f(x) = (2/5)x - 3

Therefore, the piecewise function is

f(x) = -(2/5).x -3 (2/5).r -3 -5 < x < 0 0<x<5